# Fourier growth of degree $2$ polynomials

### v1

@@ Abstract.
We prove bounds for the absolute sum of all level-$k$ Fourier coefficients for $(-1)^{p(x)}$, where polynomial $p:\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}$ is of degree $\leq 2$. First we prove the bound $\lesssim(1+\sqrt{2})^{k}$. Then, by a different method, we make it sharp: $\lesssim k^{-1/2}(1+\sqrt{2})^{k}$.
@@
LB, JS, HZ, AV:This work was completed while the authors were in residence at the Hausdorff Research Institute for Mathematics in Bonn, during the trimester program “Boolean Analysis in Computer Science”, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813; L.B. is supported by the Collaborative Research Center 1060 funded by the Deutsche Forschungsgemeinschaft and the Hausdorff Center for Mathematics, funded by the DFG under Germany’s Excellence Strategy - GZ 2047/1. O.K. was supported in part by a grant from the Israel Science Foundation (ISF Grant No. 1774/20), and by a grant from the US-Israel Binational Science Foundation and the US National Science Foundation (BSF-NSF Grant No. 2020643. J.S. is supported by Chris Umans’ Simons Institute Investigator Grant. A.V. is supported by NSF DMS-2154402 and by Hausdorff Center for Mathematics.
@@ 1. Introduction
We want to estimate independently of $n$ the $k$-th Fourier weight of $(-1)^{p(x)}$, where $p:\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}$ a polynomial of degree $\leq d$. Here, the $k$-th Fourier weight of a function $f:\mathbf{F}_{2}^{n}\to\\ mathbf{C}$ refers to the quantity
$\sum_{|A|=k}|\hat{f}(A)|\,,$
@@
$\hat{f}(A)=\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}f(x)(-1)^{x_{A}}\,,\qquad x_{A}=\sum_{i\in A}x_{i}\,.$
For polynomials $p$ of degree 00 or $1$, the absolute sum of all Fourier coefficients of $(-1)^{p(x)}$ is exactly $1$. In particular the $k$-th Fourier weight is at most $1$ for all $k$.
For polynomials of degree at most $2$, we will prove the following.
@@
For any degree $2$ polynomial $p:\mathbf{F}^{n}_{2}\to\mathbf{F}_{2}$, the $k$-th Fourier weight of $(-1)^{p(x)}$ is at most
$(1+\sqrt{2})^{k}\,.$
This confirms a conjecture of Chattopadhyay, Hatami, Hosseini and Lovett, see [CHHL, Open Problem 1.7], for polynomials of degree $2$. They state in [CHHL] that they have a proof for degree $2$ polynomomials. Pooya Hatami has kindly sketched to us what argument they had in mind, we reproduce it in Section 3. Up to a factor $k^{-1/2}$, it gives the same bound.
2. Proof of Theorem 1
We denote by $W(2,k)$ the maximum over all polynomials of degree $2$ on $\mathbf{F}_{2}^{n}$ of the $k$-th Fourier weight of $(-1)^{p(x)}$:
$W(2,k)=\max_{\begin{subarray}{c}n\geq 1\\ \mathrm{deg}\,p\leq 2\end{subarray}}\sum_{A\subset[n],|A|=k}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|\,.$ (2.1)
Here we abbreviate $x_{A}=\sum_{i\in A}x_{i}$. Theorem 1 claims that
$W(2,k)\leq(1+\sqrt{2})^{k}\,.$ (2.2)
We will prove this by induction on the number of variables $n$ and on $k$. Note that the estimate is obvious when $k\leq 0$ or $n<k$, since then we have $W(2,k)=0$. It also holds for $n=k$, since then $W(2,k)\leq 1$. So let $k\geq 1$ and $n\geq k+1$.
We denote the maximum $k$-th Fourier weight among all $n$ variable degree $2$ polynomials by $W(2,k,n)$:
$W(2,k,n)=\max_{\mathrm{deg}\,p\leq 2}\sum_{A\subset[n],|A|=k}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|\,.$
For our induction step, we assume that for all $k^{\prime}<k$
$W(2,k^{\prime})\leq(1+\sqrt{2})^{k^{\prime}}$ (2.3)
@@ 2. Proof of Theorem 1
$W(2,k,n)$ $=\sum_{A\subset[n],|A|=k}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|$
$=\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}\sum_{A\subset[n],|A|=k}\varepsilon_{A}\cdot(-1)^{x_{A}}$ (2.5)
for certain $\varepsilon_{A}\in\{-1,1\}$.
@@
$p(x)=x_{1}+\dotsc+x_{m}+const$
for some $m$. In that case (2.5) is $1$ if $k=m$ and 00 if $k\neq m$, so $W(2,k,n)\leq 1$.
If $p$ is not affine, then we can write it (perhaps after permuting the variables) as
$p(x)=p_{0}(\bar{x})+x_{1}p_{1}(\bar{x})+x_{2}p_{2}(\bar{x})+x_{1}x_{2}\,,$
where $\bar{x}=(x_{3},\dotsc,x_{n})$. For a set $A\subset[n]$, we abbreviate $\bar{A}=A\cap\{3,\dotsc,n\}=A\setminus\{1,2\}$. Writing out the sum over $x_{1},x_{2}$ in (2.5), we then obtain
$W(2,k,n)$
$=\frac{1}{4}\frac{1}{2^{n-2}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})}\cdot$
$\Big{[}\sum_{A\cap\{1,2\}=\emptyset}\varepsilon_{A}(-1)^{x_{\bar{A}}}+\sum_{A\cap\{1,2\}=\{1\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}+\sum_{A\cap\{1,2\}=\{2\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}+\sum_{\{1,2\}\subset A}\varepsilon_{A}(-1)^{x_{\bar{A}}}\Big{]}$
$+\frac{1}{4}\frac{1}{2^{n-2}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})}(-1)^{p_{1}(\bar{x})}\cdot$
$\Big{[}\sum_{A\cap\{1,2\}=\emptyset}\varepsilon_{A}(-1)^{x_{\bar{A}}}-\sum_{A\cap\{1,2\}=\{1\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}+\sum_{A\cap\{1,2\}=\{2\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}-\sum_{\{1,2\}\subset A}\varepsilon_{A}(-1)^{x_{\bar{A}}}\Big{]}$
$+\frac{1}{4}\frac{1}{2^{n-2}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})}(-1)^{p_{2}(\bar{x})}\cdot$
$\Big{[}\sum_{A\cap\{1,2\}=\emptyset}\varepsilon_{A}(-1)^{x_{\bar{A}}}+\sum_{A\cap\{1,2\}=\{1\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}-\sum_{A\cap\{1,2\}=\{2\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}-\sum_{\{1,2\}\subset A}\varepsilon_{A}(-1)^{x_{\bar{A}}}\Big{]}$
$-\frac{1}{4}\frac{1}{2^{n-2}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})}(-1)^{p_{1}(\bar{x})}(-1)^{p_{2}(\bar{x})}\cdot$
$\Big{[}\sum_{A\cap\{1,2\}=\emptyset}\varepsilon_{A}(-1)^{x_{\bar{A}}}-\sum_{A\cap\{1,2\}=\{1\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}-\sum_{A\cap\{1,2\}=\{2\}}\varepsilon_{A}(-1)^{x_{\bar{A}}}+\sum_{\{1,2\}\subset A}\varepsilon_{A}(-1)^{x_{\bar{A}}}\Big{]}.$
We will simplify this using the following identity, which is valid for all $p_{1},p_{2}\in\mathbf{F}_{2}$:
@@
Using the first identity on the first terms in square brackets, the second identity on the second terms etc. in the expansion of $W(2,k,n)$ yields that $W(2,k,n)$ equals
$\frac{1}{2}\frac{1}{2^{n-1}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})}\sum_{A\subset[n]\setminus\{1,2\},|A|=k}\varepsilon_{A}(-1)^{x_{A}}$
$+\frac{1}{2}\frac{1}{2^{n-1}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})+p_{2}(\bar{x})}\sum_{A\subset[n]\setminus\{1,2\},|A|=k-1}\varepsilon_{A\cup\{1\}}(-1)^{x_{A}}$
$+\frac{1}{2}\frac{1}{2^{n-1}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})+p_{1}(\bar{x})}\sum_{A\subset[n]\setminus\{1,2\},|A|=k-1}\varepsilon_{A\cup{2}}(-1)^{x_{A}}$
$+\frac{1}{2}\frac{1}{2^{n-1}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}-(-1)^{p_{0}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})+p_{1}(\bar{x})+p_{2}(\bar{x})}\sum_{A\subset[n]\setminus\{1,2\},|A|=k-2}\varepsilon_{A\cup\{1,2\}}(-1)^{x_{A}}\,.$
The expression in the first line, for fixed $A$, is the $A$-th Fourier coefficient of the function $(-1)^{q_{1}(\bar{x})}$, where
$q_{1}(\bar{x})=p_{0}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x}).$
This $q_{1}$ is a degree $2$ polynomial, because $p_{1}$ and $p_{2}$ are by assumption affine polynomials. Furthermore, $q_{1}$ only depends on $n-2$ variables. Hence, the absolute value of first line is at most $\frac{1}{2}W(2,k,n-2)$.
Similarly, the second and third line are in absolute value at most $\frac{1}{2}W(2,k-1,n-2)$, and the last line at most $\frac{1}{2}W(2,k-2,n-2)$. It follows that
@@
$\leq(\frac{1}{2}(1+\sqrt{2})^{k}+(1+\sqrt{2})^{k-1}+\frac{1}{2}(1+\sqrt{2})^{k-2})$

### v2

@@ Abstract.
We prove bounds for the absolute sum of all level-$k$ Fourier coefficients for $(-1)^{p(x)}$, where polynomial $p:\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}$ is of degree $1$ or degree $2$.
@@
This work was completed while the authors were in residence at the Hausdorff Research Institute for Mathematics in Bonn, during the trimester program “Boolean Analysis in Computer Science”, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813;
L.B. is supported by the Collaborative Research Center 1060 funded by the Deutsche Forschungsgemeinschaft;
J.S. is supported by Chris Umans’ Simons Institute Investigator Grant;
A.V. is supported by NSF DMS-2154402 and by the Hausdorff Center for Mathematics;
H.Z. is supported by NSF DMS-2453408.
@@ 1. Introduction
Let $\mathbf{F}_{2}=\{0,1\}.$ Any function $f:\mathbf{F}_{2}^{n}\to\mathbf{C}$ has the Fourier–Walsh expansion
$f(x)=\sum_{A\subset[n]}\widehat{f}(A)(-1)^{x_{A}},$ (1.1)
where $A\subset[n]:=\{1,2,\dots,n\}$ and $x_{A}=\sum_{i\in A}x_{i}$. The degree of $f$ is defined as $\deg(f):=\max_{\widehat{f}(A)\neq 0}|A|$, where $|A|$ denotes the cardinality of $A$.
Boolean functions $f:\mathbf{F}_{2}^{n}\to\{-1,1\}$ play an important role in theoretical computer science and social choice theory. Of particular interest to us are those of the form $f(x)=(-1)^{p(x)}$ with $p:\mathbf{F}^{n}_{2}\to\mathbf{F}_{2}$ a polynomial of small degree.
We want to estimate independently of $n$ the $k$-th Fourier weight of $(-1)^{p(x)}$, where $p:\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}$ a polynomial of degree at most $d$. Here, the $k$-th Fourier weight of a function $f:\mathbf{F}_{2}^{n}\to\mathbf{C}$ refers to the quantity
$\sum_{|A|=k}|\widehat{f}(A)|\,$,
@@
$\widehat{f}(A)=\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}f(x)(-1)^{x_{A}}\,.$
In fact, Chattopadhyay, Hatami, Hosseini and Lovett [CHHL18, Open problem 1.7] conjectured the following.
Conjecture 1.
Let $f(x)=(-1)^{p(x)},x\in\mathbf{F}_{2}^{n}$ with $p:\mathbf{F}_{2}^{n}\to\mathbf{F}_{2}$ being a polynomial of degree at most $d$. Then there exists a constant $c_{d}>0$ depending only $d$ such that
$\sum_{A:|A|=k}|\widehat{f}(A)|\leq c_{d}^{k}$
for all $k\geq 1$. In particular, the constant $c_{d}=2^{O(d)}$ should work.
As explained in [CHHL18], it is necessary to assume the exponential dependence on $k$. In case $\deg(p)\leq 1$, we may write $p(x)=a+x_{i_{1}}+\cdots+x_{i_{m}}$ for $1\leq i_{1}<\dots<i_{m}\leq n$ and some $a\in\mathbf{F}_{2}$. Then
$|\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)+x_{A}}|=\begin{cases}2^{n}&A=\{i_{1},\dots,i_{m}\}\\ 0.&\textnormal{otherwise}\end{cases}$
Therefore, we have
$\sum_{A\subset[n]}|\widehat{f}(A)|=\frac{1}{2^{n}}\sum_{A\subset[n]}|\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)+x_{A}}|=1$
and thus
$\sum_{A:|A|=k}|\widehat{f}(A)|\leq 1.$
For $p$ of degree at most 2, we will prove by a simple induction argument the following.
@@
For any polynomial $p:\mathbf{F}^{n}_{2}\to\mathbf{F}_{2}$ of degree at most $2$, the $k$-th Fourier weight of $f(x)=(-1)^{p(x)}$ is at most $(1+\sqrt{2})^{k}$, that is,
$\sum_{A:|A|=k}|\widehat{f}(A)|\leq(1+\sqrt{2})^{k}.$
This confirms a conjecture in [CHHL18] for polynomials of degree $2$. We note that the authors state in [CHHL18] to already have a proof in that case, which appears however not to be published. In Section 3, we give another proof of Theorem 1 based on the strategy they had in mind. We thank Pooya Hatami for kindly sharing it with us. It gives the following sharp upper bound.
Theorem 2.
For any polynomial $p:\mathbf{F}^{n}_{2}\to\mathbf{F}_{2}$ of degree at most $2$, the $k$-th Fourier weight of $f(x)=(-1)^{p(x)}$ is
$\sum_{A:|A|=k}|\widehat{f}(A)|\leq\sqrt{\frac{2}{\pi}}k^{-1/2}(1+\sqrt{2})^{k}.$
This is sharp in the sense that there exist pairs $(p,k)$ for which the ratio of right hand side and left hand side gets arbitrarily close to $1$.
Pseudorandom generators and fractional pseudorandom generators are constructed in [CHHL18], Section 4, given that the exponential bounds on Fourier weight of $(-1)^{p}$ for high degree polynomials $p$ are proved.
Pseudorandom generators are widely studied in complexity theory. There are several general frameworks used to approach their possible construction. The literature is vast, one can read a survey [Gol10] for example. A harmonic analysis approach was suggested in [CHHL18], and it requires the good estimate of Fourier weights. The reader can see other approaches in e. g. [NW94], [HLW06].
In [CGLLS20] it was shown that for polynomial error, Fourier bounds on the first $O(\log n)$ levels is sufficient to recover the seed length (the smallness of pseudorandom generator) obtained in [CHLT19], which required bounds on the entire tail.
2. Proof of Theorem 1 by induction
We denote by $W(2,k,n)$ the maximum over all polynomials of degree $2$ on $\mathbf{F}_{2}^{n}$ of the $k$-th Fourier weight of $(-1)^{p(x)}$:
$W(2,k,n)=\max_{\mathrm{deg}\,p\leq 2}\sum_{A\subset[n],|A|=k}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|\,.$ (2.1)
Here we abbreviate $x_{A}=\sum_{i\in A}x_{i}$. Theorem 1 claims that for all $0\leq k\leq n$
$W(2,k,n)\leq(1+\sqrt{2})^{k}\,.$ (2.2)
We will prove this by induction on the number of variables $n$ and on $k$. Note that the estimate is obvious when $k=0$ or $k=n$, since then we have $W(2,k,n)\leq 1$. So let $k\geq 1$ and $n\geq k+1$.
For our induction step, we assume that for all $k^{\prime}<k$ and all $n^{\prime}$
$W(2,k^{\prime},n^{\prime})\leq(1+\sqrt{2})^{k^{\prime}}$ (2.3)
@@ 2. Proof of Theorem 1 by induction
$W(2,k,n)$ $=\sum_{A\subset[n],|A|=k}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|.$ (2.5)
@@
$p(x)=a+x_{i_{1}}+\dotsb+x_{i_{m}}$
for some $m$ and $a\in\mathbf{F}_{2}$. In that case, as explained above, (2.5) is $1$ if $k=m$ and 0 if $k\neq m$, so $W(2,k,n)\leq 1$.
If $p$ is not affine, then we write it, perhaps after permuting the variables, as
$p(x)=p_{0}(\bar{x})+x_{1}p_{1}(\bar{x})+x_{2}p_{2}(\bar{x})+x_{1}x_{2}\,$, (2.6)
where $\bar{x}=(x_{3},\dotsc,x_{n})$. For a set $A\subset[n]$, we abbreviate $\bar{A}=A\setminus[2]=A\cap\{3,\dotsc,n\}$. According to our notation, $\bar{x}_{\bar{A}}=\sum_{i\in[n]\setminus[2]}x_{i}$. Writing out the sum over $x_{1},x_{2}$ in (2.5), we then obtain
$W(2,k,n)=I_{0}+I_{1}+I_{2}+I_{12}$
where (we omit the constraints $A\subset[n]$ and $|A|=k$ for notational convenience)
$I_{0}$ $=\sum_{A\cap\{1,2\}=\emptyset}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|,$
$I_{1}$ $=\sum_{A\cap\{1,2\}=\{1\}}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|,$
$I_{2}$ $=\sum_{A\cap\{1,2\}=\{2\}}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|,$
$I_{12}$ $=\sum_{A\cap\{1,2\}=\{1,2\}}\left|\frac{1}{2^{n}}\sum_{x\in\mathbf{F}_{2}^{n}}(-1)^{p(x)}(-1)^{x_{A}}\right|.$
Recall that by (2.6)
$p(0,0,\bar{x})=p_{0}(\bar{x}),\qquad p(0,1,\bar{x})=p_{0}(\bar{x})+p_{2}(\bar{x})$
and
$p(1,0,\bar{x})=p_{0}(\bar{x})+p_{1}(\bar{x}),\qquad p(1,1,\bar{x})=1+p_{0}(\bar{x})+p_{1}(\bar{x})+p_{2}(\bar{x}).$
Consider the term $I_{0}$ first. Since $A\cap\{1,2\}=\emptyset$, we have $(-1)^{x_{A}}=(-1)^{\bar{x}_{\bar{A}}},$ and thus
$I_{0}=\frac{1}{2^{n}}\sum_{A\cap\{1,2\}=\emptyset}\left|\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{\bar{x}_{\bar{A}}+p_{0}(\bar{x})}\left(1+(-1)^{p_{1}(\bar{x})}+(-1)^{p_{2}(\bar{x})}-(-1)^{p_{1}(\bar{x})+p_{2}(\bar{x})}\right)\right|.$
The other three terms can be rewritten in a similar way. For $I_{1}$, since $A\cap\{1,2\}=\{1\}$, we have $(-1)^{x_{A}}=(-1)^{x_{1}+\bar{x}_{\bar{A}}},$ and thus
$I_{1}=\frac{1}{2^{n}}\sum_{A\cap\{1,2\}=\{1\}}\left|\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{\bar{x}_{\bar{A}}+p_{0}(\bar{x})}\left(1-(-1)^{p_{1}(\bar{x})}+(-1)^{p_{2}(\bar{x})}+(-1)^{p_{1}(\bar{x})+p_{2}(\bar{x})}\right)\right|.$
For $I_{2}$, since $A\cap\{1,2\}=\{2\}$, we have $(-1)^{x_{A}}=(-1)^{x_{2}+\bar{x}_{\bar{A}}},$ and thus
$I_{2}=\frac{1}{2^{n}}\sum_{A\cap\{1,2\}=\{2\}}\left|\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{\bar{x}_{\bar{A}}+p_{0}(\bar{x})}\left(1+(-1)^{p_{1}(\bar{x})}-(-1)^{p_{2}(\bar{x})}+(-1)^{p_{1}(\bar{x})+p_{2}(\bar{x})}\right)\right|.$
For $I_{12}$, since $A\cap\{1,2\}=\{1,2\}$, we have $(-1)^{x_{A}}=(-1)^{x_{1}+x_{2}+\bar{x}_{\bar{A}}},$ and thus
$I_{12}=\frac{1}{2^{n}}\sum_{A\cap\{1,2\}=\{1,2\}}\left|\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{\bar{x}_{\bar{A}}+p_{0}(\bar{x})}\left(1-(-1)^{p_{1}(\bar{x})}-(-1)^{p_{2}(\bar{x})}-(-1)^{p_{1}(\bar{x})+p_{2}(\bar{x})}\right)\right|.$
We will simplify these formulae using the following Lemma.
Lemma 3.
For $a,b\in\mathbf{F}_{2}$, we have
$1+(-1)^{a}+(-1)^{b}-(-1)^{a+b}=2(-1)^{ab}\,.$ (2.7)
Proof.
The proof follows easily by direct computation. We provide here, however, a complicated proof. Note that $(-1)^{x}=1-2x,x\in\mathbf{F}_{2}$. So
$(-1)^{a+b}=(-1)^{a}(-1)^{b}=(1-2a)(1-2b)=1-2(a+b)+4ab.$
So
$1+(-1)^{a}+(-1)^{b}-2(-1)^{ab}$ $=1+1-2a+1-2b-2(1-2ab)$
$=1-2(a+b)+4ab$
$=(-1)^{a+b}.\qed$
So for polynomials $p_{1},p_{2}$ taking values in $\mathbf{F}_{2}$, we have
@@
Using these four identities, we may rewrite the above four terms as
$I_{0}=\frac{1}{2}\sum_{\bar{A}\subset[n]\setminus[2],|\bar{A}|=k}\left|\frac{1}{2^{n-2}}\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})}(-1)^{\bar{x}_{\bar{A}}}\right|,$
$I_{1}=\frac{1}{2^{n-1}}\sum_{\bar{A}\subset[n]\setminus[2],|\bar{A}|=k-1}\left|\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})+p_{2}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})}(-1)^{\bar{x}_{\bar{A}}}\right|,$
$I_{2}=\frac{1}{2^{n-1}}\sum_{\bar{A}\subset[n]\setminus[2],|\bar{A}|=k-1}\left|\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})+p_{1}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})}(-1)^{\bar{x}_{\bar{A}}}\right|,$
and
$I_{12}=\frac{1}{2^{n-1}}\sum_{\bar{A}\subset[n]\setminus[2],|\bar{A}|=k-2}\left|\sum_{\bar{x}\in\mathbf{F}_{2}^{n-2}}(-1)^{p_{0}(\bar{x})+p_{1}(\bar{x})+p_{2}(\bar{x})+p_{1}(\bar{x})p_{2}(\bar{x})}(-1)^{\bar{x}_{\bar{A}}}\right|.$
Note that the polynomial $p_{0}+p_{1}p_{2}$ on $\mathbf{F}_{2}^{n-2}$ is of degree at most $2$, so by definition,
$I_{0}\leq\frac{1}{2}W(2,k,n-2).$
Similarly, we have
$I_{1},I_{2}\leq\frac{1}{2}W(2,k-1,n-2),\quad I_{12}\leq\frac{1}{2}W(2,k-2,n-2).$
All combined, we have shown that
$W(2,k,n)\leq\frac{1}{2}W(2,k,n-2)+W(2,k-1,n-2)+\frac{1}{2}W(2,k-2,n-2).$
Therefore,
@@
$\leq\frac{1}{2}(1+\sqrt{2})^{k}+(1+\sqrt{2})^{k-1}+\frac{1}{2}(1+\sqrt{2})^{k-2}$